OpenStudy (zubhanwc3):

1) A particle is moving with velocity: v(t) = t^2 -9t+18 with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero.

OpenStudy (zubhanwc3):

\[v(t) = t^{2} - 9t + 18\]

OpenStudy (anonymous):

so what is the question?

OpenStudy (zubhanwc3):

Find each of the following. Give units in the answer. a) The initial acceleration b) The average velocity over the interval 0 to 8 seconds c) The instantaneous velocity at time 5 secs d) The time(s) when the particle is at rest e) The time interval(s) when the particle is moving left f) The time interval(s) when the particle is moving right g) The speed of the particle at time 4 secs h) The time interval(s) when the particle is (i) going faster (ii) slowing down i) Find the total distance the particle has traveled between 0 and 8 seconds

OpenStudy (zubhanwc3):

@sourwing srry, forgot to post it.

OpenStudy (anonymous):

O.O

OpenStudy (zubhanwc3):

this question is annoying me T_T

OpenStudy (anonymous):

Which part are you having trouble with?

OpenStudy (zubhanwc3):

i dont know how to find initial accel

OpenStudy (anonymous):

Acceleration is the derivative. So \(a_0 = v'(0)\)

OpenStudy (zubhanwc3):

im guessing i take the derivative, but then, what do they mean by initial.

OpenStudy (zubhanwc3):

ah. ty.

OpenStudy (zubhanwc3):

and what about average velocity?

OpenStudy (anonymous):

Given no information, we assume \(t_i=0\) and so \(a_i=a(t_i)=v'(0)\)

OpenStudy (anonymous):

Average velocity done here by using calculating the average of the function \(v\) over the interval from \(0\) to \(8\).

OpenStudy (zubhanwc3):

so i take the velocity of each, and divide by 8, or is there an easier way?

OpenStudy (anonymous):

Do you know how to calculate the average value of a function over an interval?

OpenStudy (zubhanwc3):

i forgot tbh.

OpenStudy (anonymous):

\[ \overline{f_{a\to b}} = \frac{1}{b-a}\int\limits_a^bf(t)\;dt = \frac{F(b)-F(a)}{b-a} \]Where \(F' =f\)

OpenStudy (zubhanwc3):

o, u mean the mean value theorem, so c would be the average velocity?

OpenStudy (zubhanwc3):

or in this case, f

OpenStudy (anonymous):

Welll, in terms of the mean value theorem, in this case \(c\) is the time at which \(v\) is at it's average value on the \([a,b]\) interval.

OpenStudy (zubhanwc3):

aite, ty, i should be able to handle the rest, can i message you if i get stuck again?

OpenStudy (anonymous):

Mean value theorem only tells us that at some point in our interval we reach the mean value. What I gave you resembles the mean value theorem, but its purpose is to calculate the mean value.

OpenStudy (anonymous):

If you get stuck, let me know.

OpenStudy (zubhanwc3):

ty

OpenStudy (zubhanwc3):

and what is the distance the particle has traveled.

OpenStudy (zubhanwc3):

when s(t) is positive, is it going to the left or right?

OpenStudy (anonymous):

is right positive, or is left positive?

OpenStudy (zubhanwc3):

it doesnt say, im guessing left is negative tho, so how do i find the distance tho?

OpenStudy (anonymous):

Use the following:\[ \int _0^tv(t)\;dt = s(t)-s(0) = s(t)-1 \]Assuming that positive means right.

OpenStudy (zubhanwc3):

tyvm